# Examining Input Modalities and Visual Feedback Designs in Mobile Expressive Writing

### v1

@@ 3.2. Input Modality Conditions
We asked the participants to type text with the soft keyboard in keyboard-based condition and send a voice message in voice condition. In Japan, where this study was conducted, there are different keyboard designs people commonly use besides the QWERTY layout (e.g., flick input). We allowed our participants to use any keyboard they liked to maintain the realistic setting of their technology use. For the voice message input modality, we used the voice message functionality already available in LINE, with which our study participants audio-recorded their utterances. We prohibited our participants from using speech recognition functionalities for text input as the LINE platform did not offer a way to differentiate between text generated by speech recognition or text entered with a software keyboard.
@@ 3.3.1. Independent Variables
We prepared two conditions for this part of the study: keyboard-based and voice conditions. We chose these two methods as they represent major text entry approaches commonly available in computers and smart devices (Yu et al., 2019). In using smartphones, soft keyboards are the most common text entry method while voice input is increasing its share (Enge, 2020). They also represent two common communication modalities in real-world therapy (i.e., writing down stressful experiences vs. talking to a therapist).
@@ 3.3.2. Dependent Variables for Stress-relieving Effect and Content
We captured the following four quantitative measures. For the text-related measures, we used Whisper111We decided to use Whisper because it exhibited a high level of recognition accuracy (approximately 94%) for Japanese at the time of our analysis. (Radford et al., 2022) to transcribe the collected voice data.
@@
To capture how the participants perceived the cathartic effect of each expressive writing, we asked them to respond to the question of “Comparing your responses before and after you described a stressful event, how did you feel your stress level has changed?” (Figure 1e). This question was based on previous work that captured stress level change through stress coping interventions (Tong et al., 2023). The response was in a five-point Likert scale, ranging from [-2, 2] (2: greatly increased, 1: increased, 0: no difference, -1: decreased, and -2: greatly decreased). We asked participants to respond to this question immediately after every expressive writing.
@@
To capture the post-experiment perceptions of the cathartic effect, we asked “{Keyboard-based/Voice} input was helpful for reducing my stress,” based on Howe’s work (Howe et al., 2022). The response was a five-point Likert scale (1: strongly disagree to 5: strongly agree). We collected this data in the post-experiment questionnaire.
@@
We counted the length of each expressive writing. The length was calculated as the character length of the original Japanese text, and one English word experimentally corresponds to 2–2.5 characters in Japanese in terms of word/character counts.
@@
To capture how much the participants disclosed their emotions in the expressive writing, we collected the use of words related to negative emotion. The usage of emotional words is known to relate to cathartic effect (Pennebaker and Chung, 2007). We analyzed the text with the Linguistic Inquiry and Word Count program (LIWC) (Pennebaker et al., 2001; Igarashi et al., 2022). For each participant/input modality combination, we separated the descriptions into words using MeCab (Kudo, 2005) (a widely used segmentation and morphological analysis tool for Japanese) and then fed them into LIWC. We used the score of positive/negative emotion words from the results of LIWC analysis, which is the percentage of words registered in the library as positive/negative emotion words out of the number of words that appeared in the text.
@@ 3.3.3. Dependent Variables for Usability and Preference
We collected the following three quantitative data measures in the post-experiment questionnaire to understand how the participants perceived the usability of each input modality and which modality they preferred.
@@
To capture perceptions of the ease of use, we asked a question of “It was easy to enter stress by {Keyboard/Voice},” based on previous work compared the usability of digital coping methods (Howe et al., 2022). The response was a five-point Likert scale (1: “It was difficult to do” to 5: “It was easy to do.”).
@@
To capture the perceived mental workload associated with each input modality condition, we used NASA-TLX (Hart and Staveland, 1988). NASA-TLX’s score ranges from 0 to 100, calculated as the weighted sum of the six subjective subscales.
@@
To capture the preference for input modalities, we asked a question of “Which input method would you prefer to use?” with two answering options: “keyboard-based input” and “voice input.”
@@ 3.4. Procedure
On the first day, we asked the participants to sign the consent form, fill in a pre-experiment questionnaire (see Appendix A), and ensure that they were able to use our system. The pre-experiment questionnaire included the Perceived Stress Scale (PSS) (Cohen et al., 1994) to gauge the stress levels of our participants prior to the experiment. We explained that the main objective of this study was to collect their stress-related emotion through our system instead of investigating the effect of the two input modalities. In this manner, we avoided giving participants potential biases. After the pre-experiment questionnaire, we shared the link and QR code for the installation of our chatbot. We split participants into two groups with considering the balance of gender, age, and occupations to counter-balance the order of the presentations of the two input modality conditions.
From Day 2 to Day 14, our system asked participants to share their stressful experiences and related emotions given one of the input modalities. We sent the question between 6 p.m. and 7 p.m. daily, and reminded the incomplete participants between 9 p.m. and 10 p.m. Additionally, since stress has been noted to have correlations with weekdays (Brantley et al., 1988), we aimed to minimize this effect by having participants experience two different conditions on a weekly basis. On Day 8, we flipped the input modality conditions. We set the condition lengths to 7 days and 6 days so that the first and final day, which required a relatively longer time to answer the questionnaire, occurred on weekends and made participation easier.
On the last day, we conducted a debriefing where we disclosed the true objectives of this study. We explicitly explained that they would be fully compensated even if they decided to withdraw at this point, and explicitly offered an opportunity to withdraw. None of the participants withdrew from the study. We then asked the participants to fill out a post-experiment questionnaire (Table 3 in Appendix A), which included questions about ease of use, mental workload and post-experiment cathartic effect of each input modality and the preference on input modalities, and the questions for PSS. We also collected qualitative data with the open-ended questions, asking participants for their thoughts or opinions about each input condition and their reasons for their preferences. We offered all participants approximately 19 USD in the local currency for the completion of the study.
@@ 3.5. Participant recruitment
We recruited participants through Crowdworks, one of the most widely used crowdsourcing platforms in Japan. Majima et al. (Majima et al., 2017) confirmed that the subjects available through Crowdworks are sufficiently reliable for behavioral research. We recruited 28 participants (14 female and 14 male; PA1 – PA28) in their 20s to 60s (Table 1).
@@ 4. Study 1 Results
We collected 483 expressive writing data and 479 data on immediate stress change, which recorded a response rate of 98.9%. Among 483 expressive writing data, 244 and 239 were entered under the keyboard text entry and voice message conditions, respectively. The average scores of pre- and post-experiment PSS were 28.5 ( = 7.87) and 27.6 ( = 8.95), respectively, and our Wilcoxon signed-rank test did not find a significant difference ( = 1, = 0.250, = .27, = 0.0473).
@@ 4.1. Perceived Usability & Preference
The average values of ease of use were 4.11 ( = 1.10) and 2.46 ( = 1.57) for the keyboard text entry and voice message conditions, respectively (Figure 2). Our Wilcoxon signed-rank test found a significant difference ( = 1, = -4.18, ¡ .001, = 0.79).
@@
The average values of mental workload were 51.7 ( = 16.5) and 62.0 ( = 19.0) for the keyboard input and voice message conditions, respectively (Figure 2). Our Wilcoxon signed-rank test found a significant difference ( = 1, = -3.46, ¡ .001, = 0.65). There were significant negative correlations between the ease of use and the mental workload scores in both the keyboard text entry ( = -0.39, ¡ .05) and voice message ( = -0.43, ¡ .05) conditions.
@@
Twenty-one participants preferred keyboard text entry to voice messages while seven preferred voice messages (Figure 2). Our chi-square test revealed that participants preferred the keyboard text entry significantly more compared to the voice message condition ( = 7.0, ¡ .01).
We further examined comments we obtained from our participants in the post-experiment questionnaire. We identified three major factors that highlighted the two important factors in the keyboard text entry method: privacy friendliness, and reflection through typing.
Our participants expressed that the voice message input was often inappropriate, because emotional disclosure tends to include private or sensitive information. Stressful events often occur on the job or in relationships with others, and expressive writing may include negative emotions toward others. Therefore, our participants hesitated to record voice messages in public spaces as PA1 mentioned: “I found it very challenging to use voice input because I didn’t want others to hear me, so I had to wait for moments when there were no people around. This was especially difficult when I was at work in the company.”
On the contrary, typing allowed participants to refine their thoughts and feelings because they could review and modify the content. As they typed on the keyboard, the text they entered appeared on their screen. They reviewed the content, making revisions as needed, which helped them review their writing objectively and organize their thoughts and emotions as PA12 commented “I was able to write about the events of the day while recalling them in detail, which allowed me to organize my emotions.”
In contrast, some participants said organizing the content while speaking was challenging, and they sometimes found themselves not knowing what to say in the middle of recording. PA1 commented: “I felt that I didn’t know what to say once I started voice input, and I couldn’t properly convey my thoughts.” With the keyboards, our participants were relaxed and could enter calmly: “Also, when it came to speech, I was too nervous to say what I was thinking, so texting was easier for me since I could slowly look back and express my thoughts.” (PA2).
@@ 4.2. Stress-relieving Effect & Content of Expression
The average values of immediate stress change were -0.14 ( = 0.44) and -0.085 ( = 0.50) under the keyboard input and voice message conditions, respectively (Figure 3). We ran a Mann-Whitney’s U test to evaluate the difference in the cathartic effect, but we could not find a significant effect ( = 27087, = -0.67, = .47, = 0.03).
@@
Figure 3 shows the average values of the responses to the question asking the cathartic effect (the keyboard text entry: 3.21, = 0.99; the voice message: 2.61, = 1.10). The keyboard text entry condition was rated significantly higher than the voice message condition, as indicated by our Wilcoxon signed-rank test ( = 1, = -3.85, ¡ .05, = 0.73). We also found that the post-experiment cathartic effect had a significant positive correlation with the perceived ease of use in both the keyboard text entry ( = .52, ¡ .01) and voice message conditions ( = .71, ¡ .001).
@@
The average length of emotional expression data was 65.2 characters ( = 43.8) under the keyboard condition and 63.6 ( = 38.4) under the voice message condition in Japanese 222Each data would correspond to roughly 25–32 words in English on average, for one English word experimentally corresponds to 2–2.5 characters in Japanese in terms of word/character counts.. Our Wilcoxon signed-rank test did not reveal a significant difference ( = 1, = 0.73, = 0.48, = 0.14).
@@ 5. Study 2: Examination of Effects of Visual Feedback
Our next exploration is to examine the visual feedback provided after expressive writing, related to RQ2. We obtained a separate approval for the study protocol described in this section form the Institutional Review Board of the first author’s institution.
@@ 5.1. Experiment Design
We used the same system architecture with additions of visual feedback for the expressed emotional writing. We decided to use the a keyboard-based text entry method based on our findings in the first study. We also made a minor modification to allow participants to share positive episodes when they did not have any stressful moments though we did not use such data for our later analysis.
@@ 5.1.1. Independent Variables: Visual Feedback Conditions
We added two visual feedback conditions: Passive and Interactive. The two conditions differed by the extent of embodiment (Soesilo et al., 2021) and the agency of elimination (Lee and Schwarz, 2021). These two factors might influence the outcome, referring to the previous work (Lee and Schwarz, 2021). We also included the None condition where no visual feedback was provided as a baseline, which is equivalent to the configuration used in Study 1.
@@
This condition offered participants unembodied, non-interactive visual feedback. When participants submitted their stress-related emotional expression text, the system provided a short video where a note with the submitted text exploded. As we introduced in section 2.3, materializing negative emotions and giving feedback that the object extinguished also dissolve its tied emotions (Briñol et al., 2013). In the context of mobile stress management, users may rather prefer low-burden interactions (Howe et al., 2022), and physically-demanding actions may not be suitable. This condition also represents a situation where users delegate deleting actions to others, which has not been fully explored yet (Lee and Schwarz, 2021). Figure 4 shows a walkthrough of this condition. We adopted an explosion visual effect similar to feedback used in the previous VR studies (Grieger et al., 2021; Björling et al., 2019).
@@
This condition offers participants embodied, interactive visual feedback. After submitting their stress-related emotional expression text, the system provides an interface where they can proactively explode a note with the submitted text by tapping it. This design was also based on studies on destroying materialized emotions. Figure 5 shows the walkthrough of this condition. This condition thus would require additional actions from users, but may help them build the sense of control, potentially contributing to cathartic effect positively.

### v3

@@ 3.2. Input Modality Conditions
Participants used a soft keyboard to type in the Keyboard condition and sent voice messages in the Voice condition. Conducted in Japan, the study allowed any keyboard design for realistic use, including non-QWERTY options like flick input. For voice messages, they used LINE’s existing functionality. Participants were prohibited from using speech recognition for text input since LINE could not distinguish between speech-recognized and keyboard-entered text. Whisper111We decided to use Whisper because it exhibited a high level of recognition accuracy (approximately 94%) for Japanese at the time of our analysis. (Radford et al., 2022) was used for transcribing voice data.
@@ 3.3.1. Independent Variables
We established two input conditions for the study: using soft keyboards (Keyboard) and voice message (Voice). These methods are key text entry approaches in computers and smart devices (Yu et al., 2019). While soft keyboards dominate smartphone text entry, voice input is gaining popularity (Enge, 2020). They also mirror communication styles in therapy, such as writing about stress or speaking with a therapist.
@@ 3.3.2. Dependent Variables for Stress-relieving Effect and Content
As it is known that expressive writing leads to immediate negative mood and eventual stress reduction (Esterling et al., 1994; Pennebaker and Chung, 2007), we captured stress reduction effects immediately after the intervention and in the post-experiment questionnaire.
@@
To assess the cathartic effect of expressive writing, participants answered a question about changes in stress levels before and after describing a stressful event (Figure 1e), based on previous stress coping research (Tong et al., 2023). Responses were on a five-point Likert scale from -2 (greatly decreased) to 2 (greatly increased), completed after each writing task.
@@
We assessed the cathartic effect post-experiment by asking, based on Howe (Howe et al., 2022), if “{Keyboard/Voice} input was helpful for reducing my stress,” using a five-point Likert scale (1: strongly disagree to 5: strongly agree) in the questionnaire.
@@
We determined the length of each expressive writing by counting characters in the original Japanese text, with one English word equating to 2–2.5 Japanese characters.
@@
To assess participants’ emotional disclosure in writing, we analyzed positive and negative emotional word usage with the Linguistic Inquiry and Word Count (LIWC) program (Pennebaker et al., 2001; Igarashi et al., 2022). For each participant and input modality, descriptions were word-segmented using MeCab (Kudo, 2005) and processed with LIWC. We measured the percentage of positive/negative emotion words according to LIWC results.
@@ 3.3.3. Dependent Variables for Usability and Preference
We gathered three quantitative measures in the post-experiment questionnaire to assess participants’ perceived usability and preference for each input modality.
@@
To assess ease of use, we used the question “It was easy to enter stress by {Keyboard/Voice},” aligning with earlier usability studies of digital coping methods (Howe et al., 2022). Responses were on a five-point Likert scale (1: “It was difficult to do” to 5: “It was easy to do.”).
@@
We used NASA-TLX to capture the perceived mental workload associated with each input modality condition (Hart and Staveland, 1988).
@@
To assess input modality preferences, we asked: “Which input method would you prefer, Keyboard or Voice?”
@@ 3.4. Procedure
On the first day, participants first signed the consent form, received a withdrawal of consent form, and were informed that they could discontinue the study at any time, with any collected data discarded, and would still be compensated for any completed tasks. Then they completed a pre-experiment questionnaire (see Appendix A), and confirmed their ability to use our system. The questionnaire included the Perceived Stress Scale (PSS) (Cohen et al., 1994) to assess initial stress levels. We explained that the study’s primary goal was to collect stress-related emotions, avoiding bias about input modality effects. Afterward, we provided a link and QR code to access to our system. We split participants into two balanced groups by gender, age, and occupation to counterbalance input modality presentations.
From Day 2 to Day 14, participants shared stressful experiences and emotions, but only submissions using the assigned input modality were accepted. We asked questions daily between 6 and 7 p.m., with reminders sent to incomplete participants from 9 to 10 p.m. To reduce weekday stress effects (Brantley et al., 1988), participants alternated weekly conditions. Participants who failed to respond received follow-up messages through the crowdsourcing platform. On Day 8, input modalities were switched. Conditions lasted 7 and 6 days, aligning the start and end on weekends to facilitate easier participation.
On the final day, we revealed the study’s true objectives in a debriefing. Participants were assured of full compensation even if they chose to withdraw, yet none did. They signed a new consent form and completed a post-experiment questionnaire (Table 3 in Appendix A), covering ease of use, mental workload, post-experiment cathartic effect for each input modality, and preference on input modalities, along with PSS questions. Qualitative data were also gathered through open-ended questions about their thoughts on input conditions and preferences. These data aimed to provide deeper insights into the quantitative results while maintaining participant anonymity and encouraging candid, reflective responses.
We collected the consent forms and questionnaires through Google Forms and communicated with participants via the crowdsourcing platform. We offered all participants approximately 19 USD in the local currency for the completion of the study.
@@ 3.5. Participant recruitment
We recruited participants through Crowdworks, one of the most widely used crowdsourcing platforms in Japan. Majima et al. (Majima et al., 2017) confirmed that the subjects available through Crowdworks are sufficiently reliable for behavioral research. We conducted no screening based on participants’ mental health history, as the study targeted a broad general population. We recruited 28 participants (14 female and 14 male; PA1 – 28) in their 20s to 60s (Table 1).
@@
3.6. Ethical Considerations
This study was approved by the Institutional Review Board of the first author’s institution, and all procedures complied with ethical guidelines for research involving human participants.
To avoid introducing potential biases, we employed a deception study protocol. At the beginning of the study, we explained to our participants that the objective of the study was to collect stress-related emotional expressions. At the end of the study, we disclosed the true objective of the study (i.e., a comparison of input modalities), and offered them an explicit opportunity for withdrawal without any negative consequence to them (e.g., full compensation would be offered even if they decided to withdraw). We then asked them to consent to the disclosed experimental protocol and data use. All the participants consented after this debriefing.
Before participation and after debriefing, all individuals provided informed consent through an online form. Participants were informed that they could withdraw from the study at any time without penalty. In the event of withdrawal, any data already collected would be discarded. A separate withdrawal-of-consent form was also provided, and participants were compensated for all tasks completed prior to withdrawal.
We did not collect any personally identifiable information. Participants were identified only by pseudonymous usernames assigned by the crowdsourcing platform. These were used solely for managing participation and compensation. All data were securely stored and used exclusively for this study.
To minimize potential psychological distress, we informed participants that they could stop the experiment at any time if they felt discomfort, and that we would also terminate participation if signs of excessive stress were observed. Participants were compensated fairly based on the study duration and workload.
@@ 4. Study 1 Results
In this section, we present the results of the first user study described in the previous section, focusing on usability and cathartic effects. We collected 483 expressive writing entries and 479 data points on immediate stress change, i.e. a response rate of 98.9%. Among 483 expressive writing entries, 244 and 239 were entered under the Keyboard and Voice conditions, respectively. The average scores for PSS before and after the experiment were 28.5 ( = 7.87) and 27.6 ( = 8.95), respectively, and our Wilcoxon signed-rank test did not find a significant difference ( = 1, = 0.250, = .27, = 0.0473). The total length of responses across four open-ended questions averaged approximately 306 characters per participant, ranging from 36 to 799 characters, with a response rate of 100%.
@@ 4.1. Perceived Usability & Preference
The Keyboard condition showed better usability than the Voice condition and thus was preferred. We also extracted the rationale for their preference by analyzing the qualitative data.
@@
Participants felt keyboard-based text entry was easier to use than sending Voice messages. The average values of ease of use were 4.11 ( = 1.10) and 2.46 ( = 1.57) for the Keyboard and Voice conditions, respectively (Figure 2). Our Wilcoxon signed-rank test found a significant difference ( = 1, = -4.18, ¡ .001, = 0.79).
@@
In line with the ease of use, keyboard-based text entry required less mental workload. The average values of mental workload were 51.7 ( = 16.5) and 62.0 ( = 19.0) for the Keyboard and Voice conditions, respectively (Figure 2). Our Wilcoxon signed-rank test found a significant difference ( = 1, = -3.46, ¡ .001, = 0.65). There were significant negative correlations between the ease of use and the mental workload scores in both the Keyboard ( = -0.39, ¡ .05) and Voice ( = -0.43, ¡ .05) conditions.
@@
Twenty-one participants preferred keyboard-based text entry to voice messages while seven preferred voice messages (Figure 2). Our chi-square test revealed that participants preferred the Keyboard condition significantly more compared to the Voice condition ( = 7.0, ¡ .01).
We performed qualitative analysis on the comments we obtained from our participants in the post-experiment questionnaire. We identified three major themes that highlighted the two important factors in the Keyboard text entry method: privacy friendliness, and reflection through typing.
Participants found voice message input unsuitable as it often involved private or sensitive emotional disclosures, making them hesitant to use it in public spaces, especially when expressing negative emotions about stressful events at work or in relationships, as PA1 mentioned: “I found it very challenging to use voice input because I didn’t want others to hear me, so I had to wait for moments when there were no people around. This was especially difficult when I was at work in the company.”
Typing with keyboards enabled participants to refine their thoughts and feelings by reviewing and modifying content objectively on their screens, as PA12 commented: “I was able to write about the events of the day while recalling them in detail, which allowed me to organize my emotions.”
Some participants said organizing the content while speaking was challenging, and they sometimes found themselves not knowing what to say in the middle of recording. PA1 commented: “I felt that I didn’t know what to say once I started voice input, and I couldn’t properly convey my thoughts.” With the keyboards, our participants were relaxed and could enter calmly: “Also, when it came to speech, I was too nervous to say what I was thinking, so texting was easier for me since I could slowly look back and express my thoughts.” (PA2).
@@ 4.2. Stress-relieving Effect & Content of Expression
We analyzed quantitative data on stress relief and expressive writing. Both Keyboard and Voice conditions promptly reduced stress, with participants experiencing greater relief with Keyboard. Expression content showed minimal change.
@@
The average values of immediate stress change were -0.14 ( = 0.44) and -0.085 ( = 0.50) under the Keyboard and Voice conditions, respectively (Figure 3). Note that we gained immediate stress change by calculating each participant’s average ratings on each feedback. We ran a Mann-Whitney’s U test to evaluate the difference in the cathartic effect, but we could not find a significant effect ( = 27087, = -0.67, = .47, = 0.03).
@@
Figure 3 presents the average cathartic effect ratings: Keyboard was 3.21 ( = 0.99) and Voice was 2.61 ( = 1.10). The Keyboard text entry was rated significantly higher than the Voice message, as shown by the Wilcoxon signed-rank test ( = 1, = -3.85, ¡ .05, = 0.73). A significant positive correlation existed between the cathartic effect and perceived ease of use for both Keyboard ( = .52, ¡ .01) and Voice ( = .71, ¡ .001).
@@
The average length of emotional expression data was 65.2 characters ( = 43.8) under the Keyboard condition and 63.6 ( = 38.4) under the Voice condition in Japanese 222Each data would correspond to roughly 25–32 words in English on average, for one English word experimentally corresponds to 2–2.5 characters in Japanese in terms of word/character counts.. Our Wilcoxon signed-rank test did not reveal a significant difference ( = 1, = 0.73, = 0.48, = 0.14).
@@ 5. Study 2: Examination of Effects of Visual Feedback
As mentioned earlier, to further explore design possibilities for enhancing the effects of mobile expressive writing, we conducted an experiment related to RQ2, examining the impact of visual feedback that materializes emotional catharsis. This section describes the design of our second user study. We obtained a separate approval for the study protocol described in this section form the Institutional Review Board of the first author’s institution.
@@ 5.1. Experiment Design
We used the same system architecture, adding visual feedback for emotional writing. Based on the first study, we chose a keyboard-based text entry method. Additionally, we allowed participants to share positive episodes when they did not have stressful moments although this data was not used in our analysis.
@@ 5.1.1. Independent Variables: Visual Feedback Conditions
We prepared three visual feedback conditions: None, Passive, and Interactive. The None condition offered no visual feedback, which is equivalent to the configuration used in Study 1, and serves as a baseline. For the visual feedback conditions, we grounded our design choices in prior research on materialized emotional disposal as referred to in section 2.2.2.
@@
In this condition, participants received unembodied, non-interactive visual feedback. When they submitted stress-related texts, the system showed a short video with their text on a note that exploded. Figure 4 details this process. We used an explosion effect similar to previous VR studies (Grieger et al., 2021; Björling et al., 2019). This design assumes that conveying the meaning of disposal alone can be effective without direct physical engagement (Briñol et al., 2013), and that passive interventions are particularly suitable for high-stress contexts requiring minimal effort (Zhao et al., 2023; Howe et al., 2022).
@@
In this condition, participants were offered interactive visual feedback. After submitting their stress-related text, participants could tap to explode a note with their text (Figure 5 shows the walkthrough). This design draws on grounded cognition theory (Barsalou, 2008), suggesting that physical engagement strengthens the perceived meaning of actions and supports emotional de-fusion (Lee and Schwarz, 2021; Hatvany et al., 2018). It also promotes a sense of agency (Moore, 2016), potentially enhancing cathartic effects.
By comparing these conditions, we aimed to examine how varying levels of user engagement in emotional disposal influence the psychological outcomes in mobile expressive writing.